# A comprehensive exploration of quasisymmetric stellarators and their coil sets

### v2

@@ 5.6 Quasiaxisymmetry
For completeness, in Figure 14, we complete a PCA on quasiaxisymmetric devices in QUASR with $n_{\text{fp}}=2$ and $\iota=0.4$. The cumulative PC ratio is just over $80\%$, revealing that a non-negligible amount of information is lost in this projection, and it might be useful to include a third principal component in the visualization. Given the two principal axes, we project devices outside QUASR onto the figure and label them. This area of the parameter space is well explored. In Figure 15, we do a PCA for QA devices, subselected based on their rotational transform ($\iota=0.5,0.6$) and number of field periods ($n_{\text{fp}}=2,3$). Labeled devices from the literature appear similar to the devices in QUASR. We do not provide cross sections of hypothetical devices on the linear manifold as we did in previous sections because, visually, the cross sections do not differ significantly from each other.
@@ 6 Discussion and conclusion
The main goal of this paper was to introduce a new data set of optimized stellarators, and provide tools to navigate the contents of it. We start by highlighting a few devices in QUASR first, and then move to analysing the data-set. First, we frame QUASR within the near-axis framework and its predicted landscape of quasisymmetric configurations. We find close correspondence of devices in QUASR to regions of favorable near-axis quasisymmetry. However, this approach does not apply to the full set of devices in QUASR given the difference in assumptions that went into their construction, most notably constraints on elongation and simplicity of the shaping. A more model agnostic procedure to analyse the data is then presented in the form of a PCA. On the subsets of the data set that we have explored in this work, we find that the first and second principal components act predominantly on the elongation (and different field twisting), and relative positions of cross sections of the magnetic surfaces. These two dimensions are able to capture the devices in the data set to a large degree. For some subsets of the data, we observe clear clusters. Within the clusters, we find evidence that the devices form a continuum whose shape follows that of the design constraint manifold.
PCA is not only a very useful dimensionality reduction and visualization tool, it can also be used to create compact representations of the data that can be used for initial stages of optimization. For the sets presented, little information is lost when using two to three principal components. This is true of even more subsets throughout the QUASR database, as illustrated in Figure 16, where we compute a PCA for various combinations of $n_{\text{fp}}$ and $\iota$ for QA and QH devices. It is possible to attain a cumulative PC ratio of 0.9 consistently. This illustrates that much of the variation between the devices can be characterized on a low-dimensional, linear manifold. This points to the lower dimensional nature of the subset of quasisymmetric configurations in the larger space of stellarators. Whether the information lost in the dimensionality reduction is significant depends on the application. For the purposes of visualization in this work, a cumulative PC ratio of 90% appeared satisfactory. But in other applications, this might not be the case.
@@
This is useful because very simple, linear formulas can be used to generate warm starts for optimization. Alternatively, one might attempt optimizing in the PC subspace. These directions are not explored in this paper.
@@ Appendix C Reference QS configurations from the literature
BULLER_0p50 2 6.00 0.50 (m) QA 15 Buller et al. (2024)
BULLER_0p60 2 6.00 0.60 (m) QA 15 Buller et al. (2024)
NIES_AR=8 3 8.00 0.42 (e) QA 15 Nies et al. (2024)
NIES_AR=6p5 3 6.50 0.42 (e) QA 15 Nies et al. (2024)
@@
To make a direct comparison of these to the devices in the QUASR dataset, we must construct their corresponding feature vectors using the following procedure. First, we take the geometry of the device’s outermost surface and use VMEC (Hirshman & Whitson, 1983b) to determine the vacuum magnetic field on the entire toroidal volume. We take the surface in the device that has aspect ratio closest to 10 for QA and 12 for QH. The Boozer coordinates on this surface are computed using BOOZ_XFORM (Sanchez et al., 2000). Finally, we compute the Fourier harmonics of this surface parametrized in Boozer coordinates. We may then treat them as we did with QUASR objects, and when necessary project them to the appropriate principal component plots. We make direct comparison of these configurations to the relevant ones in QUASR when they share the value of $n_{\text{fp}}$, helicity, and mean rotational transform (out to aspect ratio 10 (for QA) and 12 (for QH) to an accuracy of ±$\pm 0.05$). In that case, we may represent them alongside each other, as we do in Figures 69, 10 or 14.

### v3

@@ 5.6 Quasiaxisymmetry
For completeness, on the left of Figure 14, we complete a PCA on quasiaxisymmetric devices in QUASR with $n_{\text{fp}}=2$ and $\iota=0.4$. The cumulative PC ratio is just over $80\%$, revealing that a non-negligible amount of information is lost in this projection, and it might be useful to include a third principal component in the visualization. Given the two principal axes, we project devices outside QUASR onto the figure and label them. This area of the parameter space is well explored. In Figure 14, we do a PCA for QA devices, subselected based on their rotational transform ($\iota=0.6$) and number of field periods ($n_{\text{fp}}=2,3$). Labeled devices from the literature appear similar to the devices in QUASR. We do not provide cross sections of hypothetical devices on the linear manifold as we did in previous sections because, visually, the cross sections do not differ significantly from each other.
@@ 6 Discussion and conclusion
The main goal of this paper was to introduce a new data set of optimized stellarators, and provide tools to navigate the contents of it. We start by highlighting a few devices in QUASR first, and then move to analysing the data-set. First, we frame QUASR within the near-axis framework and its predicted landscape of quasisymmetric configurations. We find close correspondence of devices in QUASR to regions of favorable near-axis quasisymmetry. However, this approach does not apply to the full set of devices in QUASR given the difference in assumptions that went into their construction, most notably constraints on elongation and simplicity of the shaping. A more model agnostic procedure to analyze the data is then presented in the form of a PCA. On the subsets of the data set that we have explored in this work, we find that the first and second principal components act predominantly on the elongation (and different field twisting), and relative positions of cross sections of the magnetic surfaces. These two dimensions are able to capture the devices in the data set to a large degree. For some subsets of the data, we observe clear clusters. Within the clusters, we find evidence that the devices form a continuum whose shape follows that of the design constraint manifold.
For the sets presented, two to three principal components appear sufficient for visualization. This is true of even more subsets throughout the QUASR database, as illustrated in Figure 15, where we compute a PCA for various combinations of $n_{\text{fp}}$ and $\iota$ for QA and QH devices. It is possible to attain a cumulative PC ratio of 0.9 consistently. This illustrates that much of the variation between the devices can be characterized on a low-dimensional, linear manifold. This points to the lower dimensional nature of the subset of quasisymmetric configurations in the larger space of stellarators. This echoes the analysis in (Sengupta et al., 2023), where it was revealed that quasisymmetry imposes a lower-dimensionality on the field strength. We would like to highlight though that this lower-dimensionality does not necessarily extend to the coil design space, which is known to be complex with multiple minima (Giuliani, 2024; Wechsung et al., 2022a). This is because there are many, vastly different coil sets that produce approximately the same magnetic field (Landreman, 2017). This highlights that some sort of global search or a problem reformulation is indeed necessary for coil design.
Whether the information lost in the dimensionality reduction is significant depends on the application. For the purposes of visualization in this work, a cumulative PC ratio of 90% appeared satisfactory. But in other applications, this might not be the case.
@@
Some useful extensions of this include using the very simple, linear formulas learned during the PCA to generate warm starts for optimization. Alternatively, one might attempt optimizing in the PC subspace. These directions are not explored in this paper.
@@ Appendix C Reference QS configurations from the literature
BULLER_0p60 2 6.00 0.60 (m) QA 14 Buller et al. (2024)
NIES_AR=6p5 3 6.50 0.42 (e) QA 14 Nies et al. (2024)
@@
To make a direct comparison of these to the devices in the QUASR dataset, we must construct their corresponding feature vectors using the following procedure. First, we take the geometry of the device’s outermost surface and use VMEC (Hirshman & Whitson, 1983b) to determine the vacuum magnetic field on the entire toroidal volume. We take the surface in the device that has aspect ratio closest to 10 for QA and 12 for QH. The Boozer coordinates on this surface are computed using BOOZ_XFORM (Sanchez et al., 2000). Finally, we compute the Fourier harmonics of this surface parametrized in Boozer coordinates. We may then treat them as we did with QUASR objects, and when necessary project them to the appropriate principal component plots. We directly compare these configurations to the relevant ones in QUASR when they share the value of $n_{\text{fp}}$, helicity, and mean rotational transform (out to aspect ratio 10 (for QA) and 12 (for QH) to an accuracy of ±$\pm 0.05$).